OpenStudy (princess123):

Find three consecutive odd integers such that the product of the first and second exceeds the third by 8. the first integer is ? the second integer is ? the third integer is ?

OpenStudy (anonymous):

(a + 2) * (a + 1) = a + 8 a² + 3a + 2 = a + 8 a² + 2a - 6 = 0

OpenStudy (princess123):

sorry im confused... so whats the first integer>